# PTQ4VM: Post-Training Quantization for Visual Mamba

### v1

@@ 6.1 Quantization Setting
PTQ4VM (ours) 83.7 83.4 82.2
Table 2: ImgaeNet Top-1 validation accuracy comparison of quantization methods on various models. LocalVim indicates a model that uses the CLS token.
@@ 6.2 Image Classification
The experimental results for the Vim family with CLS tokens reveal significant accuracy degradation when using MinMax, even at W8A8, as it fails to account for token-wise variance. Similarly, SmoothQuant struggles to retain essential CLS token information due to the same issue, leading to a 29.6% accuracy loss at 6-bit for Vim-B. In contrast, PTQ4VM addresses all the challenges posed by Visual Mamba, demonstrating strong performance with less than 0.5% accuracy loss at 8-bit for Vim-T/S and LocalVim, while maintaining acceptable quality at 4-bit, where other models fall short. Despite Vim-B having higher FP16 accuracy than Vim-S, its quantized performance lags across all methods, primarily due to outliers exceeding magnitudes of 20K. These findings suggest that Visual Mamba models with excessively large activation magnitudes may be inherently vulnerable to compression techniques.
In the VMamba family, all methods show lossless quality at W8A8 compared to FP16, largely due to its significantly smaller activation ranges compared to other model families. This unique characteristic can be linked to the absence of gating functions, as seen in Fig. 2b, which differentiates it from other architectures. However, while other methods struggle with W4A4, only PTQ4VM manages to maintain quality on VMamba-S/B with less than 0.4% degradation, achieving lossless quality compared to FP16 models.
@@ 6.3 Object Detection and Instance Segmentation
To demonstrate that PTQ4VM performs in downstream tasks, we evaluated its performance in Object Detection and Instance Segmentation tasks. We used the official checkpoint trained with Mask R-CNN[11] 3X MS schedule using VMamba-T backbone for our experiments, evaluating on images cropped to 1280$\times$800. As reported in Table 3, while MinMax and SmoothQuant both fail at W4A4, PTQ4VM demonstrates superior performance with less than 1.9% score degradation. At W6A6, SmoothQuant shows lower scores than MinMax, which can be attributed to the degradation caused by the difficulty imposed on weights by smoothing. In contrast, PTQ4VM achieves near-lossless performance with less than 0.3% APb and 0.1% APm degradation, due to optimal tuning by JLSS. Notably, the qualitative results in Fig. 7 validate that PTQ4VM maintains the network’s quality well, preserving even small details.
@@
Vim-T Per-tensor static 37.86 32.58 1.16$\times$
@@ 6.5 Ablation Study
For VMamba, although there is no CLS token, accounting for token-wise variance still enhances accuracy at 6-bit and 4-bit levels. Additionally, by using a grid search to optimize the truncation level and minimize L2 loss at the layer level (labeled with +Truncation), both Vim and VMamba show significant accuracy improvements—approximately 40% and 50%, respectively, at 4-bit. This highlights the importance of proper truncation in static quantization. JLSS, designed to further optimize truncation, yields additional gains of 1.7% and 10% at 6-bit and 4-bit, respectively, while boosting accuracy by 10% in VMamba-T.
@@ A.1 Quantization Configurations
In this section, we describe the detailed quantization configuration for reproducing PTQ4VM. In the Image Classification task, we trained for 10 epochs for W8A8, 50 epochs for W6A6, and 100 epochs for W4A4. The learning rates used in our experiments are reported in Table 7. We denoted the learning rate for the smoothing scale $s$ as lr_s and the learning rate for the quantization parameters (step size $\Delta_{X}$ and $\Delta_{W}$) as lr_q. Additionally, all models used in the experiments used pretrained checkpoints provided by the official repository of Vision Mamba [40], VMamba [22], and LocalMamba [12].
Vim-Ti Vim-S Vim-B VMamba-T/S/B
lr_s 1e-2 1e-3 1e-2 1e-4
lr_q 5e-4 5e-4 5e-4 1e-4
LocalVim-T LocalVim-T LocalVim-S LocalVMamba-S
lr_s 1e-2 1e-2 1e-4 1e-4
lr_q 5e-4 5e-4 5e-4 1e-4
Table 7: Learning rates for Image Classification task.

### v2

@@ 6.1 Quantization Setting
PTQ4VM (ours) 83.7 83.4 81.2
Table 2: ImageNet Top-1 validation accuracy comparison of quantization methods on various models. LocalVim indicates a model that uses the CLS token.
@@ 6.2 Image Classification
The experimental results for the Vim family with CLS tokens reveal significant accuracy degradation when using MinMax, even at W8A8, as it fails to account for token-wise variance. Similarly, SmoothQuant struggles to retain essential CLS token information due to the same issue, leading to a 29.6% accuracy loss at 6-bit for Vim-B. In contrast, PTQ4VM addresses all the challenges posed by Visual Mamba, demonstrating strong performance with less than 0.5% accuracy loss at 8-bit for Vim-Ti/S and LocalVim-T, while maintaining acceptable quality at 4-bit, where other methods fall short. Despite Vim-B having higher FP16 accuracy than Vim-S, its quantized performance lags across all methods, primarily due to outliers exceeding magnitudes of 20K. These findings suggest that Visual Mamba models with excessively large activation magnitudes may be inherently vulnerable to compression techniques.
In the VMamba family, all methods show lossless quality at W8A8 compared to FP16, largely due to its significantly smaller activation ranges compared to other model families. This unique characteristic can be linked to the absence of gating functions, as seen in Fig. 2b, which differentiates it from other architectures. However, while other methods struggle with W4A4, only PTQ4VM manages to maintain quality on VMamba-S/B with less than 1.3% degradation.
@@ 6.3 Object Detection and Instance Segmentation
To demonstrate that PTQ4VM performs in downstream tasks, we evaluated its performance in Object Detection and Instance Segmentation tasks. We used the official checkpoint trained with Mask R-CNN[11] 3X MS schedule using VMamba-T backbone for our experiments, evaluating on images cropped to 1280$\times$800. As reported in Table 3, while MinMax and SmoothQuant both fail at W4A4, PTQ4VM demonstrates superior performance with less than 3.5% score degradation. At W6A6, SmoothQuant shows lower scores than MinMax, which can be attributed to the degradation caused by the difficulty imposed on weights by smoothing. In contrast, PTQ4VM achieves near-lossless performance with less than 0.3% APb and 0.2% APm degradation, due to optimal tuning by JLSS. Notably, the qualitative results in Fig. 7 validate that PTQ4VM maintains the network’s quality well, preserving even small details.
@@
Vim-Ti Per-tensor static 37.86 32.58 1.16$\times$
@@ 6.5 Ablation Study
For VMamba, although there is no CLS token, accounting for token-wise variance still enhances accuracy at 6-bit and 4-bit levels. Additionally, by using a grid search to optimize the truncation level and minimize L2 loss at the layer level (labeled with +Truncation), both Vim and VMamba show significant accuracy improvements—approximately 40% and 50%, respectively, at 4-bit. This highlights the importance of proper truncation in static quantization. JLSS, designed to further optimize truncation, yields additional gains of 10.0% and 7.3% at Vim and VMamba, respectively.
@@ A.1 Quantization Configurations
In this section, we describe the detailed quantization configuration for reproducing PTQ4VM. The hyperparameters used in our image classification experiments are reported in Table 7. We denoted the learning rate for the smoothing scale $s$ as lrs and the learning rate for the quantization parameters (step size $\Delta_{X}$ and $\Delta_{W}$) as lrq. Following SmoothQuant [36], we introduce the exponent value $\alpha$ used to initialize the smoothing scale as a hyperparameter. Here, $\alpha$ corresponds to the exponent term in Eq. 17. In the case of Eq. 14, please note that $\alpha$ is set to 0.5.
$\mathbf{s}=\frac{\max(|\mathbf{X}|)^{\alpha}}{\max(|\mathbf{W}|)^{1-\alpha}}\in\mathbb{R}^{D_{in}}.$ (17)
All models used in the experiments used pretrained checkpoints provided by the official repository of Vision Mamba [40], VMamba [22], and LocalMamba [12].
Bits Vim-Ti Vim-S Vim-B VMamba-T VMamba-S
lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch
8-bit 1e-2 5e-4 0.5 10 1e-3 5e-4 0.5 10 1e-2 5e-4 0.5 10 1e-4 1e-4 0.5 10 1e-4 1e-4 0.5 10
6-bit 50 50 50 1e-5 1e-5 0.5 50 1e-5 1e-5 0.5 50
4-bit 100 100 100 1e-4 1e-4 0.65 1e-4 1e-4 0.6
Bits LocalVim-T LocalVim-T LocalVim-S VMamba-B LocalVMamba-S
lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch lrs lrq $\alpha$ epoch
8-bit 1e-2 5e-4 0.5 10 1e-2 5e-4 0.5 10 1e-2 5e-4 0.5 10 1e-4 1e-4 0.5 10 1e-4 1e-4 0.5 10
6-bit 50 50 50 0.5 50 0.5 50
4-bit 100 100 100 0.65 0.65
Table 7: Hyperparameters of different models for image classification across quantization settings.